# UAT-Lite: Inference-Time Uncertainty-Aware Attention for Pretrained Transformers

### v1

@@ 5.2 Baselines, Metrics, and Training Overview
We evaluate our approach against BERT-base, MC Dropout, Temperature Scaling (TS), Deep Ensembles, and representative Bayesian last-layer methods. All models share the same pretrained backbone, fine-tuning recipe, and evaluation protocol; detailed baseline configurations are described in App.˜C.
Calibration is evaluated using Expected Calibration Error (ECE), alongside robustness under distribution shift and selective prediction metrics. Results are reported as mean standard deviation over five random seeds, with deterministic baselines exhibiting negligible variance. The proposed method introduces no additional trainable parameters and operates entirely at inference time.
@@
Unless otherwise stated, we use a single uncertainty penalty , set once on development data and held fixed across all tasks and datasets. §˜D.1 shows that calibration performance is stable across a broad range of and Monte Carlo budgets , indicating that the method does not rely on sensitive hyperparameter tuning in practice across the evaluated settings.
@@ 6.1 General NLP Calibration Performance
Table 1 reports multi-seed calibration results (mean standard deviation over five seeds), computed using identical ECE binning and confidence definitions across methods. Deterministic baselines exhibit negligible variance, while stochastic methods show small but consistent variability.
UAT-Lite improves calibration relative to the unscaled fine-tuned BERT baseline, reducing average ECE from 0.117 to 0.094 (19.9% relative reduction). The largest gain is observed on MNLI (0.110 0.067), a task characterized by semantic ambiguity, where uncertainty-aware attention is particularly beneficial. These improvements persist relative to both uniform and component-specific MC Dropout, indicating that the gains arise from uncertainty-weighted attention rather than from stochasticity introduced by sampling.
Temperature scaling (TS) achieves the lowest in-domain ECE across all tasks, as expected for a post-hoc method explicitly optimized to minimize calibration error on held-out validation data. Our objective is complementary rather than competitive: while TS rescales output probabilities, it does not alter the internal representations or attention patterns that produce a prediction. In contrast, UAT-Lite introduces uncertainty-aware internal computation at inference time, allowing epistemic uncertainty to influence token interactions during contextualization. Because TS operates purely at the output level, it can be directly stacked on top of UAT-Lite without modification; we focus here on isolating the effects of uncertainty-aware attention and evaluate its benefits under distribution shift and selective prediction in subsequent sections.
Among Bayesian-style alternatives, SNGP achieves lower ECE but requires architectural modification and specialized training. Deep Ensembles provide a strong reference point (Avg ECE 0.089), but incur substantial training and storage overhead due to maintaining multiple model replicas. In contrast, UAT-Lite operates on a single pretrained transformer and introduces uncertainty handling entirely at inference time. As a result, it substantially narrows the calibration gap to Deep Ensembles while preserving deployment efficiency.
| Method | SQuAD (ECE ) | MNLI (ECE ) | SST-2 (ECE ) | Avg (ECE ) |
| BERT-base | 0.186 0.000 | 0.110 0.000 | 0.055 0.000 | 0.117 |
| MC Dropout (uniform ) | 0.173 0.001 | 0.090 0.001 | 0.039 0.001 | 0.101 |
| MC Dropout (component-specific) | 0.186 0.000 | 0.109 0.001 | 0.054 0.001 | 0.116 |
| Temp. Scaling | 0.057 0.000 | 0.028 0.000 | 0.029 0.000 | 0.038 |
| SNGP | 0.001 0.000 | 0.021 0.000 | 0.009 0.000 | 0.010 |
| VI-LastLayer | 0.119 0.001 | 0.096 0.000 | 0.136 0.002 | 0.117 |
| Deep Ensemble (5) | 0.168 0.000 | 0.060 0.000 | 0.037 0.000 | 0.089 |
| UAT-Lite | 0.163 0.002 | 0.067 0.001 | 0.051 0.004 | 0.094 |
| vs. Baseline | 12.5% | 39.0% | 6.6% | 19.9% |
@@ 6.2 Distribution Shift Robustness
We evaluate calibration robustness under distribution shift on MNLI by transferring models fine-tuned on the matched (in-domain) split to the mismatched (out-of-distribution) split. Results in Table 2 are reported as mean standard deviation over five seeds, with ECE computed consistently.
Absolute ECE values differ from the in-domain setting because MNLI shift evaluation is framed as a binary decision (entailment vs. non-entailment), and confidence is computed from the Monte Carlo–averaged predictive distribution using the maximum softmax probability. This affects the scale of ECE but preserves meaningful relative comparisons across methods.
BERT-base exhibits clear calibration drift under shift, with ECE increasing from the matched to the mismatched split (). In contrast, UAT-Lite yields lower ECE on both in-domain and out-of-domain splits and reduces calibration drift (), resulting in a substantially lower average ECE across the two conditions.
Importantly, this evaluation isolates the effect of uncertainty-aware internal computation under shift. Post-hoc calibration methods such as temperature scaling are intentionally omitted here, as they rescale output confidences without modifying the internal representations that give rise to distributional sensitivity. The improved stability observed for UAT-Lite therefore reflects the benefit of allowing epistemic uncertainty to modulate attention during contextualization, rather than relying solely on output-level correction. Overall, these results indicate that uncertainty-weighted attention produces more stable confidence estimates under matched-to-mismatched distribution shift, a setting where overconfident errors are particularly costly.
| Method | ID ECE | OOD ECE | ECE | Avg ECE |
| BERT-base | 0.2893 0.0000 | 0.3017 0.0000 | 0.0124 | 0.2955 |
| UAT-Lite | 0.0219 0.0007 | 0.0145 0.0007 | -0.0074 | 0.0182 |
@@ 6.3 Selective Prediction Performance
Table 3 reports selective prediction performance using AURC (lower is better), averaged over five seeds. All selection criteria are determined on validation data and held fixed across methods. Because the Baseline is deterministic, its variance across seeds is negligible, whereas UAT-Lite exhibits small but consistent variability across runs.
Selective prediction directly evaluates the quality of uncertainty estimates rather than calibration alone: a reliable model should abstain on difficult inputs while maintaining high coverage on confident predictions. Accordingly, this evaluation isolates whether uncertainty estimates meaningfully align with prediction risk under identical settings.
On MNLI, UAT-Lite achieves substantially higher selective coverage, increasing from 0.714 to 0.869, while reducing AURC from 0.164 to 0.053. This reflects a markedly improved risk–coverage trade-off, where abstention is applied preferentially to higher-risk inputs rather than uniformly.
On SQuAD 2.0, coverage increases from 0.501 to 0.690 and AURC decreases from 0.447 to 0.289, indicating more effective abstention behavior on ambiguous or unanswerable inputs. These gains demonstrate that uncertainty-weighted attention yields confidence estimates that better reflect underlying uncertainty.
Post-hoc calibration methods such as temperature scaling are intentionally omitted from this analysis, as they modify output probabilities without altering the confidence ranking that governs abstention behavior. These improvements reflect uncertainty-aware internal computation.
| Dataset | Method | Coverage@0.9 | Coverage@0.8 | Coverage@0.7 | AURC |
| MNLI | Baseline | 0.7139 0.0000 | 0.7527 0.0000 | 0.7926 0.0000 | 0.1640 0.0000 |
| MNLI | UAT-Lite | 0.8692 0.0003 | 0.9030 0.0006 | 0.9275 0.0010 | 0.0529 0.0002 |
| SQuAD 2.0 | Baseline | 0.5007 0.0000 | 0.5071 0.0000 | 0.5050 0.0000 | 0.4472 0.0000 |
| SQuAD 2.0 | UAT-Lite | 0.6904 0.0006 | 0.7047 0.0009 | 0.7163 0.0010 | 0.2892 0.0007 |
6.4 Computational Efficiency
Table 4 reports inference cost and calibration performance on SQuAD 2.0 under a single-accelerator, sequential inference setting. We focus on methods with clearly defined deployment behavior and directly measurable inference overhead. Predictive accuracy is not reported in this analysis, as all compared methods exhibit comparable task performance on SQuAD 2.0 and the purpose of this section is to isolate calibration behavior under differing inference-time costs.
UAT-Lite with incurs higher latency than the deterministic baseline (86.31 ms vs. 17.69 ms), reflecting the cost of repeated stochastic forward passes, and yields a throughput of 13.9 examples/sec. Under the same sequential evaluation protocol, this latency is comparable to that of a 5-model Deep Ensemble (88.80 ms), while requiring only a single pretrained model (109M parameters) rather than storing and maintaining multiple independently trained replicas (547M parameters).
Temperature scaling adds negligible runtime overhead relative to the baseline, but operates purely as an output-level post-hoc calibration method and does not alter internal computation. Overall, UAT-Lite incurs additional inference-time cost in exchange for improved calibration, while avoiding the substantial training and storage overhead associated with ensemble-based approaches and preserving a single-model deployment footprint effectively.
| Method | Params | Samples | Latency (ms) | Memory (GB) | Throughput | ECE |
| BERT-base | 109M | 1 | 17.69 | 1.78 | 68.9 | 0.186 |
| Temperature Scaling | 109M | 1 | 17.71 | 1.78 | 69.0 | 0.057 |
| Deep Ensemble (5) | 547M | 5 | 88.80 | — | — | 0.168 |
| UAT-Lite | 109M | 10 | 86.31 | 1.81 | 13.9 | 0.163 |
6.5 Negation-Aware Attention
Fig.˜3 qualitatively illustrates how uncertainty-weighted attention modulates token interactions under linguistic negation. We compare attention patterns from a standard transformer and UAT-Lite under negated linguistic constructions.
While absolute attention maps remain similar due to softmax normalization, the difference heatmap reveals small but structured redistributions. Uncertainty-weighted attention slightly downweights sentiment-bearing tokens (e.g., “positively”) while reallocating attention toward negation cues and predicate structure associated with the phrase “did not respond” in context.
These changes are intentionally modest and should be interpreted as directional rather than large-scale attention reconfiguration. Accordingly, this example is illustrative rather than a precise token-level attribution. Corpus-level analysis shows that similar directional trends recur across negated and uncertainty-sensitive sentences (see §˜D.4 for additional qualitative evidence).
@@ 6.6 Ablation Study
Fig.˜4 decomposes UAT-Lite into its constituent components to isolate their individual contributions. Uncertainty-weighted attention alone reduces ECE from 0.169 to 0.134 (21.1% relative reduction), indicating that attention modulation based on epistemic uncertainty is the primary driver of calibration improvement under controlled comparisons. Embedding-level uncertainty alone slightly degrades calibration, suggesting that stochasticity without uncertainty-aware computation is insufficient. In contrast, combining uncertainty-weighted attention with confidence-aware decision shaping yields the largest improvement (28.2% relative reduction), exceeding either component in isolation and demonstrating their complementarity.
@@ 6.7 Generalization Across BERT Models
The largest gains occur for mid-sized and domain-specialized models, including BERT-base on SQuAD, BioBERT on PubMedQA, and ClinicalBERT on MedQA, where epistemic uncertainty arising from semantic ambiguity is prominent. Improvements are smaller for models with strong baseline calibration (e.g., RoBERTa-base on MNLI) and diminish or reverse for very small or highly overparameterized models.
These results indicate that uncertainty-aware attention generalizes reliably within the BERT encoder family. Extending the mechanism to fundamentally different architectures, such as decoder-only or state-space models, remains an important direction for future work.
@@
Additional analyses on adversarial robustness, linguistic uncertainty, clinical-domain transfer, and layer-wise uncertainty diagnostics are provided in the App.˜D.
@@ 7 Conclusion
We introduced UAT-Lite, an inference-time framework that integrates epistemic uncertainty into transformer computation through uncertainty-weighted attention and a layer-wise variance diagnostic. Across NLP benchmarks, UAT-Lite reduces Expected Calibration Error relative to unscaled fine-tuned BERT (e.g., average ECE 0.117 0.094) and improves selective prediction behavior, including higher coverage at fixed confidence thresholds and lower AURC. The method also exhibits stable calibration under distribution shift in MNLI matchedmismatched transfer. While inference-time Monte Carlo sampling incurs additional latency, the approach avoids the training and storage costs of ensemble-based alternatives and requires no modification of pretrained weights. Overall, UAT-Lite provides a complementary mechanism for uncertainty-aware transformer inference, bridging the gap between post-hoc calibration and Bayesian modeling for transformers.
@@
Our framework incurs additional computational cost due to Monte Carlo sampling, with a latency increase at . This overhead is appropriate for offline, batch, or document-level applications where confidence reliability is prioritized but is prohibitive for latency-critical settings. In practical deployments, this trade-off suggests that uncertainty-aware inference should be selectively enabled for high-risk or ambiguous inputs, rather than applied uniformly across all predictions. Accordingly, UAT-Lite is best viewed as an intermediate solution between lightweight post-hoc calibration and more expensive ensemble-based approaches, rather than a universal replacement.
Finally, on clinical benchmarks (MedQA and PubMedQA), calibration gains are modest, reflecting dataset properties such as factual emphasis and limited ambiguity; these experiments are therefore treated as domain-transfer stress tests rather than clinical validation.

### v2

@@ 5.2 Baselines, Metrics, and Training Overview
We compare our approach against BERT-base, MC Dropout, TS, Deep Ensembles, and representative Bayesian last-layer methods. All models share the same pretrained backbone, fine-tuning recipe, and evaluation protocol; detailed baseline configurations are provided in App.˜C. Calibration is evaluated using Expected Calibration Error (ECE), together with robustness under distribution shift and selective prediction metrics. Results are reported as mean standard deviation over five random seeds, while deterministic baselines exhibit negligible variance. The proposed method introduces no additional trainable parameters and operates entirely at inference time.
Negation diagnostic.
We include a controlled negation perturbation as a diagnostic of uncertainty responsiveness under semantic destabilization. Because negation may alter the true label, we emphasize label-free stability signals (e.g., Margin and Entropy) rather than treating this setting as a robustness test of correctness. Full details and examples are provided in §˜D.9.
@@
Unless otherwise stated, we use a single uncertainty penalty , which is selected once on development data and then held fixed across all tasks and datasets. §˜D.11 shows that calibration performance remains stable across a broad range of values and Monte Carlo budgets , indicating that the method does not rely on sensitive hyperparameter tuning in practice for the evaluated settings.
In particular, across and , ECE varies within a narrow band (range ), while accuracy changes are negligible. Increasing improves calibration, with diminishing returns beyond , suggesting that stable behavior does not require expensive Monte Carlo budgets or fine-grained tuning.
@@ 6.1 General NLP Calibration Performance
Table 1 reports multi-seed calibration results on SQuAD, MNLI, and SST-2 (ECE ; mean standard deviation over five seeds) under a shared evaluation protocol. Across tasks, UAT-Lite improves calibration relative to the unscaled fine-tuned BERT baseline (e.g., average ECE decreases from to ), with the largest improvement on MNLI (). It also outperforms MC Dropout baselines, including uniform and component-specific variants, suggesting that the gains are not explained by stochasticity alone.
Fig. 4 shows that global- or document-level MC Dropout improves over the deterministic baseline but remains behind TS, which achieves the lowest in-domain ECE. We therefore view UAT-Lite as complementary to TS: UAT-Lite modulates internal evidence aggregation, whereas TS rescales output logits. The bottom panel of Table 1 reports MedQA and PubMedQA (single seed; accuracy reported in separate columns) as a clinical domain-transfer stress test.
MedQA analysis by linguistic bucket (negation, numeric expressions, biomedical terminology, and hedging markers; Table 12) shows a consistent pattern: UAT-Lite improves accuracy across all phenomena, while UAT-Lite +TS attains the lowest calibration error in every bucket. Overall, TS is strongest for in-domain probability calibration, UAT-Lite provides complementary uncertainty-aware inference, and UAT-Lite +TS yields the best trade-off in our evaluations.
Panel A: General NLP (multi-seed) Method SQuAD (ECE ) MNLI (ECE ) SST-2 (ECE ) Avg (ECE ) BERT-base 0.1868 0.000 0.0816 0.000 0.0531 0.000 0.1072 Base+TS 0.0788 0.000 0.0148 0.000 0.0163 0.000 0.0366 MC Dropout (uniform ) 0.173 0.001 0.090 0.001 0.039 0.001 0.101 MC Dropout (component-specific) 0.186 0.000 0.109 0.001 0.054 0.001 0.116 SNGP 0.001 0.000 0.021 0.000 0.009 0.000 0.010 VI-LastLayer 0.119 0.001 0.096 0.000 0.136 0.002 0.117 Deep Ensemble (5) 0.168 0.000 0.060 0.000 0.037 0.000 0.089 Global MC Dropout (M=10) 0.1600 0.000 0.0565 0.000 0.0501 0.000 0.0889 UAT-Lite (M=10) 0.1647 0.000 0.0638 0.000 0.0608 0.000 0.0964 UAT-Lite+TS (M=10) 0.0760 0.000 0.0201 0.000 0.0191 0.000 0.0384 Panel B: Clinical QA (single seed) Method MedQA (ECE ) MedQA (Acc ) PubMedQA (ECE ) PubMedQA (Acc ) Baseline 0.033 0.246 0.114 0.620 Base+TS 0.023 0.246 0.128 0.620 Deep Ensemble (5) 0.007 0.268 0.051 0.520 UAT-Lite 0.040 0.249 0.103 0.640 UAT-Lite+TS 0.0069 0.251 0.1018 0.630
@@ 6.2 Distribution Shift Robustness
We evaluate uncertainty under covariate shift in two settings: (i) MNLI transfer (matchedmismatched; mild shift) and (ii) four OOD suites with stronger shifts (HANS, ANLI, SNLI, IMDb). Table 2 reports MNLI transfer results (mean standard deviation over five seeds), and Table 5 reports selective prediction on OOD suites, including admitted counts .
Calibration baselines and composability with TS.
Temperature scaling (TS) is the strongest post-hoc baseline for reducing marginal ECE via global logit rescaling. By contrast, UAT-Lite modulates attention-level evidence aggregation using MC-dropout uncertainty and composes naturally with TS. On MNLI transfer, UAT-Lite +TS matches TS-level average ECE with essentially unchanged accuracy (Table 2).
Selective prediction at fixed confidence thresholds.
We report Cov@, Acc@, , and AURC for (Table 5). Across OOD suites, TS reliably reduces ECE but does not uniformly dominate selective prediction, especially when score compression reduces high- coverage. Empirically, UAT-Lite +TS is the most reliable configuration, preserving TS-level calibration while retaining attention-level gating.
Shortcut-stress test on HANS.
HANS probes heuristic shortcut reliance. We additionally report a per-heuristic breakdown and a non-entailment-only subset (Table 6), which isolates the dominant failure mode of overconfident non-entailment errors. On this diagnostic, UAT-Lite and UAT-Lite +TS improve accuracy over Base and Base+TS, suggesting that uncertainty-weighted attention mitigates shortcut-driven evidence beyond output-only rescaling.
| Method | ID ECE | OOD ECE | ECE | Avg ECE | ID Acc | OOD Acc |
| Base | 0.08090.0040 | 0.05550.0000 | 0.0040 | 0.06820.0020 | 0.71620.0048 | 0.74590.0000 |
| Base+TS | 0.01470.0031 | 0.02080.0022 | 0.00610.0040 | 0.01770.0018 | 0.71620.0048 | 0.74590.0000 |
| Global MC Dropout (M=10) | 0.05860.0037 | 0.03340.0023 | 0.0033 | 0.04600.0026 | 0.70900.0047 | 0.73980.0012 |
| UAT-LITE (M=10) | 0.06500.0028 | 0.04020.0012 | 0.0038 | 0.05260.0010 | 0.70990.0038 | 0.73950.0014 |
| UAT-LITE+TS (M=10) | 0.01630.0022 | 0.01990.0013 | 0.00350.0017 | 0.01810.0016 | 0.71050.0047 | 0.73940.0004 |
@@ 6.3 Selective Prediction Performance
Table 4 reports threshold-based selective prediction results on MNLI and SQuAD 2.0, including TS and global MC Dropout baselines. Confidence is defined as . For , we report coverage (Cov@), admitted accuracy (Acc@), admitted count (), and AURC (risk–coverage; ). Thresholds are selected on validation data for each method and then applied at test time; we also report to make coverage shifts explicit.
Selective prediction evaluates whether uncertainty aligns with prediction risk: a reliable model should abstain on higher-risk inputs while retaining coverage on lower-risk, high-confidence examples. This setting is particularly informative under fixed confidence thresholds, where improved marginal calibration does not necessarily imply better abstention behavior.
Effect of TS under fixed thresholds.
We include TS because, although it preserves the predicted label, probability rescaling changes which examples exceed a fixed threshold . As a result, TS can substantially alter threshold-level behavior even when accuracy remains unchanged. On SQuAD 2.0, for example, TS can markedly reduce Cov@0.9, reflecting confidence compression rather than a reporting artifact.
Effect of uncertainty-aware inference.
By contrast, UAT-Lite modifies internal evidence aggregation through uncertainty-aware attention. This changes how confidence is formed before thresholding and can improve alignment between confidence and risk at fixed operating points. Overall, UAT-Lite, and especially UAT-Lite +TS, yield more reliable risk–coverage behavior, whereas TS alone can substantially shift coverage through score compression; see Table 4 and Table 5.
6.4 Computational Efficiency and Latency
Because UAT-Lite relies on Monte Carlo sampling, it incurs additional inference-time cost that scales approximately linearly with the MC budget under a sequential sampling loop (i.e., stochastic forward passes). To quantify end-to-end overhead in a single canonical setting, we report wall-clock latency on an NVIDIA A100 GPU with batch size and sequence length . UAT-Lite uses stochastic passes implemented as a sequential Monte Carlo loop; under this setting, the deterministic model requires ms per batch, whereas UAT-Lite requires ms, corresponding to a overhead. Temperature scaling adds negligible test-time cost ( ms) because it only applies a scalar rescaling of the logits. Table 18 reports the resulting overhead. At , UAT-Lite incurs a substantial slowdown relative to deterministic inference, whereas TS remains essentially free. Accordingly, UAT-Lite is best suited to offline or batch evaluation, or to selectively triggered regimes (e.g., invoked only for low-confidence or high-risk inputs), rather than as an always-on mechanism in latency-critical deployments.
@@ 6.5 Ablation Study
Fig.˜5 decomposes UAT-Lite into its constituent components to isolate their individual contributions. Uncertainty-weighted attention alone reduces ECE under controlled comparisons, indicating that attention modulation based on epistemic uncertainty is a major driver of the observed calibration gains. By contrast, embedding-level stochasticity alone can slightly degrade calibration, suggesting that stochasticity without structured uncertainty-aware computation is insufficient. Combining uncertainty-weighted attention with confidence-aware decision shaping yields the largest improvement among the evaluated components, consistent with their complementarity.
To examine the design choice of where to apply uncertainty modulation within attention, we evaluate variants that gate the query (Q), key (K), value (V), or all three jointly (QKV). Fig.˜3 summarizes the calibration–accuracy trade-offs across SST-2, MNLI, and SQuAD 2.0. Overall, K-only achieves the lowest ECE on SST-2 and MNLI but can substantially reduce accuracy on SQuAD 2.0, suggesting overly aggressive evidence gating in span-style question answering. Q-only provides the most balanced behavior, improving calibration while largely preserving accuracy across tasks, whereas V-only best preserves, and in some cases slightly improves, accuracy while yielding more modest ECE gains. Joint QKV gating can further reduce ECE on some datasets but tends to over-regularize and degrade accuracy.
@@ 6.6 Generalization Across BERT Models
The largest gains are observed for mid-sized and domain-specialized models, including BERT-base on SQuAD, BioBERT on PubMedQA, and ClinicalBERT on MedQA, where epistemic uncertainty associated with semantic ambiguity appears more prominent. By contrast, improvements are smaller for models with stronger baseline calibration (e.g., RoBERTa-base on MNLI) and diminish, or occasionally reverse, for very small or heavily overparameterized models.
These results suggest that uncertainty-aware attention generalizes reliably within the BERT encoder family. Extending this mechanism to fundamentally different architectures, such as decoder-only or state-space models, remains an important direction for future work.
@@
Additional analyses are provided in Apps.˜D, E and F, including threshold-based selective prediction, OOD evaluations (e.g., HANS), global and document-level MC dropout diagnostics, component ablations, penalty-form ablations, aggregation and normalization diagnostics, scale and geometry diagnostics for token uncertainty, negation diagnostics with qualitative attention illustration, behavioral analysis by linguistic bucket on MedQA, sensitivity analyses for and , character-level adversarial robustness, full linguistic uncertainty analysis, cross-model and scaling analyses, and implementation, complexity, latency, and reproducibility details.
@@ 7 Conclusion
We introduced UAT-Lite, an inference-time framework that integrates epistemic uncertainty into transformer inference through uncertainty-weighted attention and a layer-wise variance diagnostic. Across general NLP benchmarks, UAT-Lite improves calibration relative to an unscaled fine-tuned BERT baseline and yields more reliable selective prediction behavior under fixed confidence thresholds. Under distribution shift (MNLI matchedmismatched and OOD evaluation suites), uncertainty-aware attention provides benefits that are complementary to output-level calibration: temperature scaling remains the strongest post-hoc baseline for ECE reduction, whereas UAT-Lite targets internal evidence aggregation and composes cleanly with TS when calibrated probabilities are required. Although Monte Carlo sampling increases inference latency, UAT-Lite requires neither retraining nor modification of pretrained weights, providing a practical bridge between lightweight post-hoc calibration methods and higher-cost Bayesian or ensemble-based alternatives.
@@
Compute and latency.
UAT-Lite introduces additional inference cost due to Monte Carlo dropout sampling. Under a single canonical end-to-end GPU benchmark (batch size , sequence length , sequential MC loop on one GPU), the overhead is substantial (e.g., approximately at ), whereas temperature scaling adds negligible cost (approximately ). As a result, UAT-Lite is not intended for latency-critical, always-on deployment. Instead, it is better suited to offline or batch settings, or to selectively triggered use (e.g., enabling uncertainty-aware attention only for low-confidence or high-risk inputs), where improved confidence reliability and internal evidence control may justify the additional runtime. Overall, UAT-Lite occupies an intermediate point between lightweight post-hoc calibration methods (e.g., TS) and higher-cost uncertainty methods (e.g., deep ensembles), trading additional compute for internal uncertainty-aware control signals and diagnostic capability.
Compute-once token uncertainty (no per-layer recomputation).
For efficiency and conceptual clarity, we compute token-level uncertainty once at the embedding stage (via stochastic embedding samples) and reuse this signal to modulate attention throughout the network; deeper layers are analyzed through variance attribution rather than by re-estimating after each mixing operation. Recomputing token uncertainty at intermediate layers is a potentially useful extension, but it would require defining token-level uncertainty after contextual mixing (e.g., residual connections and attention aggregation) and would further increase inference cost due to additional stochastic passes. We leave per-layer recomputation, and its associated cost–benefit trade-offs, to future work, especially in settings where additional compute is acceptable.
Not a replacement for TS in-domain.
Because temperature scaling is a strong lightweight calibrator for in-domain ECE, UAT-Lite should not be viewed as a universal substitute for TS. Our primary contribution is different: uncertainty-guided modulation of the internal evidence aggregation process through attention at inference time. Accordingly, we emphasize compositional use: UAT-Lite can be combined with TS (by fitting on development data using MC-mean logits) when the goal is to retain internal uncertainty-aware behavior while further improving output calibration.
Clinical transfer is a stress test, not clinical validation.
On clinical QA benchmarks (MedQA and PubMedQA), calibration gains can be modest, which may reflect dataset properties (e.g., a stronger emphasis on factual recall and less ambiguity in evidence aggregation) as well as evaluation constraints. We therefore treat these experiments as domain-transfer stress tests rather than evidence of clinical readiness. Any real-world clinical deployment would require task- and setting-specific validation, including broader coverage, robustness checks, and appropriate governance safeguards.
@@
This supplementary material provides additional experimental results, methodological details, and reproducibility information supporting the main paper.
Appendix A presents the mathematical foundations of UAT-Lite, including the diagnostic layer-wise variance decomposition, its derivation, and the Bayesian interpretation of MC dropout.
Appendix B provides the inference-time algorithm for uncertainty-aware attention.
Appendix C details baselines, evaluation metrics (ECE, robustness metrics, selective prediction), and reporting conventions (seeds, fairness, confidence definitions).
Appendix D reports additional experimental analyses:
Appendix E evaluates generalization across BERT-family backbones and model sizes, including cross-domain calibration and scaling effects.
Appendix F documents implementation details, compute/latency overhead, hardware setup, and code availability for reproducibility.
Appendix G provides an extended related work discussion.
Appendix H provides an AI usage statement.